OpenStudy (anonymous):

A sample of oxygen gas has a volume of 3.27 L at 27°C and 400.0 torr. How many oxygen molecules does it contain?

OpenStudy (aaronq):

Find moles with PV=nRT Convert moles to molecules with: \(n=\dfrac{N}{N_A}\) N=molecules, \(N_A\)=avogadro's number, n=moles

OpenStudy (anonymous):

n = PV / RT = (400.0 torr) × (3.27 L) / ((62.36367 L Torr/K mol) × (27 + 273)K) = 0.06991 mol (0.06991 mol ) × (6.022 × 10^23 molecules/mol) = 4.21 × 10^22 molecules O2